OpenStudy (anonymous):

state the domain and range for the following relation.... {(x,y):y=1/2x^2-5}

OpenStudy (anonymous):

\[y=\frac{1}{2x^2-5}\] set \[2x^2-*5=0\] \[x^2=\frac{5}{2}\] \[x=\pm\sqrt{\frac{5}{2}}\] so all numbers except those two

OpenStudy (anonymous):

range from \[(-\infty,-\frac{1}{5}]\cup (0,\infty)\]